OpenStudy (anonymous):

I need help! I'll Give You A Medal (: Evaluate -6 - (-4). A. -10 B. -2 C. 24 D. 2

OpenStudy (anonymous):

-6 + 4 =

OpenStudy (anonymous):

-2?...

OpenStudy (igreen):

-6 - (-4) = -6 + 4

OpenStudy (igreen):

Yep, -2 is our answer.

OpenStudy (jdoe0001):

-6 - (-4) => -6 -1 * (-4).